Four ball players have $233 between them. Melvin has $20 more than Harvey, $53 more than Bruce and $71 more than Rollo. How much money did each player have?
Answer by mananth(16946)

Melwin has x
Harvey x-20
Bruce x-53
Rollo x-71
Add them up total is 233
4x-144=233
4x= 144+233
4x=377
/4
x= 94.25
Amount in Dollars
Harvey = 94.25-20= 74.25
Bruce = 94.25-53= 41.25
Rollo= 94.25-71=23.25
Melwin = 94.25